OpenStudy (anonymous):

which is smallest number that can be subtracted from 1936v so that on being divided by 9, 10, 15, the remainder is 7 every time

ganeshie8 (ganeshie8):

you need to solve below :- 1936-x = 9p + 7 1936-x = 10q + 7 1936-x = 15r + 7

ganeshie8 (ganeshie8):

combining first two equations, gives u :- 1936-7-x = 90m 1929 - x = 90m that gives u : x = 39

ganeshie8 (ganeshie8):

since x = 39 is satisfied by 3rd equation also, we're done. 39 is the required smallest number

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yw